The wavelength of a radio wave varies inversely as its frequency. A wave with a frequency of 900 kilohertz has a length of 400 meters. What is the length of a wave with a frequency of 450 kilohertz?
Answer by richwmiller(17219)

w=k/f
900=k/400
k=360000
w=k/f
w=360000/450
w=800